# Christian Röver, Tim Friede

### v1

@@ 3.1.2 Variations of the MAP prior
The heterogeneity ($\tau$) prior was specified as a $\mathrm{half\mbox{-}Normal}(0.5)$ distribution, which is a reasonably conservative choice for endpoints such as HRs, as it covers “reasonable” and up to “fairly high” levels of heterogeneity ($\tau\leq 1$) and leaves a small prior probability for “fairly extreme” amounts ($\tau>1$) 26, 9. Since conclusions heavily depend on the heterogeneity prior settings, it may however be interesting to investigate the effects of alternative specifications; in particular, we will consider different prior scales and different distribution families.
@@
Half-normal distributions are a common and obvious choice as heterogeneity priors, possible reasons may be familiarity and availability, as well as a “flat” shape near the origin and a rather short tail 9. Variations of the distribution family commonly do not alter conclusions dramatically as long as they cover a similar range, as manifested e.g. in a common prior median.
@@
Figure 3: Illustration of MAP-prior’s dependence on heterogeneity prior distribution family. The different heterogeneity priors shown here all share the same prior median
@@
One can see that heavier-tailed heterogeneity priors also yield correspondingly heavier-tailed MAP-priors. The heterogeneity and the corresponding MAP priors’ properties are also summarized and compared in Table 2 in terms of prior quantiles and effective sample sizes ($\mathrm{ESS}_{\mathrm{ELIR}}$).
@@ 3.2 Heart failure example
The Spirit-HF trial has been designed in order to test the efficacy of spironolactone in patients with heart failure (HF) 27. Spironolactone is expected to reduce cardiovascular mortality as well as hospitalizations due to heart failure, and had previously been investigated in the Topcat trial 28. Both studies refer to the composite of (recurrent) HF hospitalization and cardiovascular death as the primary endpoint to evaluate treatment efficacy. Despite a sizeable sample size of 3445 patients and a mean follow-up duration of more than $3$ years, the Topcat trial failed to demonstrate statistical significance; the estimated hazard ratio (HR) was at 0.89 (0.77, 1.04) ($p=0.14$) 28.
The analysis of the new Spirit-HF trial meanwhile is being planned, and may take into consideration the evidence already generated in the Topcat trial. One idea may be to derive a shrinkage estimate, anticipating some between-study heterogeneity, and dynamically borrowing information from the earlier study based on the corresponding MAP prior 6, 7. For the between-study heterogeneity $\tau$, use of a $\mathrm{half\mbox{-}Normal}(0.25)$ prior may be appropriate. The heterogeneity prior may be motivated referring to anticipated levels of heterogeneity based on general considerations 9 or using empirical evidence, in particular in view of the similar study designs and the effect measure being a log-HR 29.
In the present case, analysis is based on the logarithmic HR; the HR estimated in the Topcat trial corresponds to a log-HR of $-0.117$ with a standard error of $0.077$. The corresponding Topcat likelihood along with the resulting MAP prior is illustrated in Figure 6.
@@
The variance (squared standard error) of the Topcat study’s estimate was $s_{1}^{2}=0.077^{2}$ while for the assumed heterogeneity prior the expected heterogeneity variance is $\mathrm{E}[\tau^{2}]=0.25^{2}$ (see also Table 3), so that the resulting MAP prior’s variance (6) is $s_{1}^{2}+2\,\mathrm{E}[\tau^{2}]=0.362^{2}$, and the majority of the variance is due to epistemic uncertainty relating to the anticipated similarity of the Topcat and Spirit-HF estimates. The 95% prediction interval for the MAP prior is centered at the Topcat log-HR estimate and ranges from $-0.899$ to $+0.665$, corresponding to HRs in the range [$0.407$, $1.945$]. According to the MAP prior, the probability of a beneficial treatment effect (a log-HR below zero) is 71%. The MAP prior has an ESSELIR of 399, i.e, only 12% of the 3445 actual Topcat patients, and 31% of the estimated enrolment of 1300 Spirit-HF patients 27.
@@ 4 Discussion
Despite the seemingly odd notion of a meta-analysis of a single study, the use of MAP priors remains completely consistent down to the extreme case of only one data point. The “usual” toolbox remains available, including common prior specifications 9, computation of effective sample sizes (ESS) 20 robustification 5, as well as common meta-analysis software (e.g., the bayesmeta or RBesT R packages) 4, 30 for practical implementation. In addition, for $k=1$, there are connections to bias allowance and power prior models (see Section 2) that may help motivating such an approach. MAP priors based on few estimates are generally rather heavy-tailed, which will ensure robust operating characteristics 19, 6, 7. For few data points in general, and in particular for only a single data point, the prior specification for the heterogeneity parameter $\tau$ gains in importance and needs to be particularly well-founded and convincing 9.
Empirical MAP priors may then be utilized in different ways, either to simply motivate a reasonable sample size (or other design aspects), or to implement explicit borrowing of historical information 31, 32. When MAP priors are used to also inform the analysis, it is important to approach the evaluation of operating characteristics from a sensible angle 33, 34, 35.
Van Zwet et al. (2024) argue that also the analysis of a single study should account for heterogeneity of the treatment effect across studies. Therefore, they propose to consider analyses of individual studies also within an overarching NNHM framework similar to our approach presented here; using informative, empirically motivated priors for both $\mu$ and $\tau$, inference may then be focused on the overall mean effect ($\mu$) rather than the study-specific $\theta_{1}$ even in the analysis of only a single study 36.
@@ Appendix A Appendix
A.1 Heterogeneity priors
Figure 7 shows several prior densities discussed in Section 3.1.2 that are all scaled to a common median (of 0.34, the median of a HN(0.5) distribution).
Figure 7: Illustration of the several heterogeneity priors compared in Section 3.1.2 in terms of their probability density functions. All priors are scaled such that they have a common median of 0.34 (the median of a half-normal(0.5) prior; dashed line)
Figure 8: Illustration of the several heterogeneity priors compared in Section 3.1.2 in terms of their cumulative distribution functions (CDFs). All priors are scaled such that they have a common median of 0.34 (dashed line)
A.2 Variance of the MAP prior
@@
Table 3: Expected values of $\tau^{2}$ based on various common (prior) distributions for $\tau$, depending on their scale parameter $s$. An asterisk ($\ast$) indicates that there is no simple analytical expression. The half-Cauchy prior would be an additional option, but does not have a finite expectation (it is in fact also a special case of the half-Student-$t$ prior, with $\nu=1$ degree of freedom)
@@
$\tau\sim\mathrm{half}\mbox{-}\mathrm{Student}\mbox{-}t_{\nu}(s)$ast $\ast$ $2\sqrt{\frac{\nu}{\pi}}\frac{\Gamma(\frac{\nu+1}{2})}{\Gamma(\frac{\nu}{2})(\nu-1)}s$ $(\nu>1)$ $\frac{\nu}{\nu-2}s^{2}$ $(\nu>2)$
@@
Table 3 below summarizes expectations for $\tau^{2}$ for a range of common heterogeneity priors. Note also the related Table B1 in the appendix of Röver et al. (2021) 9 giving some additional details for the prior distributions shown here.
A.3 Power prior exponent’s distribution
For any fixed heterogeneity value $\tau$, the (conditional) MAP prior is equivalent to a power prior with exponent $a_{0}=\bigl{(}2\frac{\tau^{2}}{s_{1}^{2}}+1\bigr{)}^{-1}$ (see also Section 2.5) 6. Through this functional relationship, any prior density $p_{\star}(\tau)$ for the heterogeneity implies a corresponding prior for the exponent $a_{0}$ with probability density function
@@ A.3 Power prior exponent’s distribution
Figure 9 illustrates such densities for an example value of $s_{1}=0.451$ (as in the Alport example from Section 3.1).
Figure 9: Illustration of prior distributions for the power prior exponent $a_{0}$ corresponding to certain prior distributions assumed for the heterogeneity $\tau$ (and $s_{1}=0.451$)
@@
A.4 Example R code
A.4.1 Alport example
@@
A.4.2 Heart failure example

### v2

@@ 3.1.2 Variations of the MAP prior
The heterogeneity ($\tau$) prior was specified as a $\mathrm{half\mbox{-}Normal}(0.5)$ distribution, which is a reasonably conservative choice for endpoints such as HRs, as it covers “reasonable” and up to “fairly high” levels of heterogeneity ($\tau\leq 1$) and leaves a small prior probability for “fairly extreme” amounts ($\tau>1$) 31, 12. Since conclusions heavily depend on the heterogeneity prior settings, it may however be interesting to investigate the effects of a range of reasonable alternative specifications; in particular, we will consider different prior scales and different distribution families. Among the various assumptions implemented in the analysis, a “too optimistic” heterogeneity prior (favouring small heterogenity) might yield results inappropriately close to a common-effect analysis, while overly “pessimistic” or “conservative” assumptions may on the other hand may eventually lead to very little borrowing of information.
@@
Half-normal distributions are a common and obvious choice as heterogeneity priors, possible reasons may be familiarity and availability, as well as a “flat” shape near the origin and a rather short tail 12. Variations of the distribution family commonly do not alter conclusions dramatically as long as they cover a similar range, as manifested e.g. in a common prior median.
@@
Figure 3: Illustration of MAP-prior’s dependence on the heterogeneity prior distribution family. The different heterogeneity priors shown here all share the same prior median
@@
One can see that heavier-tailed heterogeneity priors also yield correspondingly heavier-tailed MAP-priors.
@@
The heterogeneity and the corresponding MAP priors’ properties are also summarized and compared in Table 2 in terms of prior quantiles and effective sample sizes ($\mathrm{ESS}_{\mathrm{ELIR}}$).
@@
It is sometimes also instructive to observe the effects of variations of the prior on the resulting estimates; e.g., varying the heterogeneity prior scale allows for a sensitivity (or tipping point) analysis. Such an analysis is shown in Appendix A.6; the amount of borrowing in reflected in the shrinkage interval’s width, but in the present example inference would not change qualitatively, and a log-HR of zero always remains included.
@@ 3.2 Heart failure example
The Spirit-HF trial has been designed in order to test the efficacy of spironolactone in patients with heart failure (HF) 32. Spironolactone is expected to reduce cardiovascular mortality as well as hospitalizations due to heart failure, and had previously been investigated in the Topcat trial 33. Both studies refer to the composite of (recurrent) HF hospitalization and cardiovascular death as the primary endpoint to evaluate treatment efficacy. Despite a sizeable sample size of 3445 patients and a mean follow-up duration of more than $3$ years, the Topcat trial failed to demonstrate statistical significance; the estimated hazard ratio (HR) was at 0.89 (0.77, 1.04) ($p=0.14$) 33.
The analysis of the new Spirit-HF trial meanwhile is being planned, and may take into consideration the evidence already generated in the Topcat trial. One idea may be to derive a shrinkage estimate, anticipating some between-study heterogeneity, and dynamically borrowing information from the earlier study based on the corresponding MAP prior 9, 10. For the between-study heterogeneity $\tau$, use of a $\mathrm{half\mbox{-}Normal}(0.25)$ prior may be appropriate. The heterogeneity prior may be motivated referring to anticipated levels of heterogeneity based on general considerations 12 or using empirical evidence, in particular in view of the similar study designs and the effect measure being a log-HR 34.
In the present case, analysis is based on the logarithmic HR; the HR estimated in the Topcat trial corresponds to a log-HR of 0.117-0.117 with a standard error of $0.077$. The corresponding Topcat likelihood along with the resulting MAP prior is illustrated in Figure 6.
@@
The variance (squared standard error) of the Topcat study’s estimate was $s_{1}^{2}=0.077^{2}$ while for the assumed heterogeneity prior the expected heterogeneity variance is $\mathrm{E}[\tau^{2}]=0.25^{2}$ (see also Table A3), so that the resulting MAP prior’s variance (6) is $s_{1}^{2}+2\,\mathrm{E}[\tau^{2}]=0.362^{2}$, and the majority of the variance is due to epistemic uncertainty relating to the anticipated similarity of the Topcat and Spirit-HF parameters. The 95% prediction interval for the MAP prior is centered at the Topcat log-HR estimate and ranges from 0.899-0.899 to +0.665+0.665, corresponding to HRs in the range [$0.407$, $1.945$]. According to the MAP prior, the probability of a beneficial treatment effect (a log-HR below zero) is 71%. The MAP prior has an $\mathrm{ESS}_{\mathrm{ELIR}}$ of 399, i.e, only 12% of the 3445 actual Topcat patients, and 31% of the estimated enrolment of 1300 Spirit-HF patients 32. This means that the prior derived from the Topcat data will not enter the eventual analysis as an additional 3445 patients (as would be the case if both study populations were pooled naïvely), but instead we expect an accuracy corresponding to a total of some $1300+399$ patients. The Spirit-HF study’s contribution to its own shrinkage estimate may also be assessed 10; assuming that both studies show the same dependence of standard error and sample size, the Spirit-HF study will account for a minimum of $61\%$ in weight to the eventual effect estimate.
@@ 4 Discussion
Despite the seemingly odd notion of a meta-analysis of a single study, the use of MAP priors remains completely consistent down to the extreme case of only one data point. The “usual” toolbox remains available, including common prior specifications 12, computation of effective sample sizes (ESS) 23 robustification 8, as well as common meta-analysis software (e.g., the bayesmeta or RBesT R packages) 7, 35 for practical implementation. In addition, for $k=1$, there are connections to bias allowance and power prior models (see Section 2) that may help help motivating a MAP approach (or vice versa). MAP priors based on few estimates are generally rather heavy-tailed, which will ensure robust operating characteristics 22, 9, 10. For few data points in general, and in particular for only a single data point, the prior specification for the heterogeneity parameter $\tau$ gains in importance and needs to be particularly well-founded and convincing 12.
While only the normal model (NNHM) was discussed here, the idea also extends to other model families; for example, derivation of a MAP prior would also work for a binomial-normal model (as implemented in the RBesT package) 35. Another related approach (with some similarity to the power prior) is given by the commensurate prior 36, which, however does not constitute a special case of the MAP prior. Empirical MAP priors may then be utilized in different ways, either to simply motivate a reasonable sample size (or other design aspects), or to implement explicit borrowing of historical information 37, 38. When MAP priors are used to also inform the analysis, it is important to approach the evaluation of operating characteristics from a sensible angle; the naive application of classically “frequentist” measures to judge a Bayesian procedure, in particular when informative priors are involved, will often not provide a meaningful assessment of its actual features 39, 40, 41.
A common concern in the context of the use of historical data is that an informative prior might unduly dominate the eventual analysis; for example, in the heart failure example application, one might be worried that the much larger Topcat trial would swamp the data from the smaller Spirit-HF study. However, for the shrinkage estimate of interest here, the second study’s contribution is bounded by a minimum of $61\%$ within the suggested setup. This proportion would increase for a more conservative heterogeneity prior specification 12 or when implementing robustification 8, however, such modelling decisions should probably rather be based on considerations of prior information than on deduced operating characteristics.
Besides considerations of the value of “borrowed” information for a given parameter estimate, MAP priors based on historical data may also be interesting for the design of subsequent trials, with or without the eventual use of shrinkage estimation in the final analysis. Historical information may then help determining sensible ranges for nuisance parameters 37 or sample sizes 42, 43, for interim decisions 8, 44, or it may be used in a more comprehensive fashion to to ensure a positive joint outcome 44, 45.
Van Zwet et al. (2024) argue that also the analysis of a single study should account for heterogeneity of the treatment effect across studies. Therefore, they propose to consider analyses of individual studies also within an overarching NNHM framework similar to our approach presented here; using informative, empirically motivated priors for both $\mu$ and $\tau$, inference may then be focused on the overall mean effect ($\mu$) rather than the study-specific $\theta_{1}$ even in the analysis of only a single study 46, 47.
@@ Appendix A Appendix
A.1 Posterior predictive distribution
When an informative $\mathrm{Normal}(\mu_{\mathrm{p}},\sigma_{\mathrm{p}}^{2})$ prior is assumed for the overall mean $\mu$, the posterior predictive distribution for a “new” study-specific mean $\theta_{k+1}$ (conditional on a given heterogeneity value) is again normal with moments
$\mathrm{E}[\theta_{k+1}|y_{1},\ldots,y_{k},s_{1},\ldots,s_{k},\tau]$ $=$ $\frac{\frac{\mu_{\mathrm{p}}}{\sigma_{\mathrm{p}}^{2}}+\sum_{i=1}^{k}\frac{y_{i}}{s_{i}^{2}+\tau^{2}}}{\frac{1}{\sigma_{\mathrm{p}}^{2}}+\sum_{i=1}^{k}\frac{1}{s_{i}^{2}+\tau^{2}}}$ (9)
$\mathrm{Var}(\theta_{k+1}|y_{1},\ldots,y_{k},s_{1},\ldots,s_{k},\tau)$ $=$ $\frac{1}{\frac{1}{\sigma_{\mathrm{p}}^{2}}+\sum_{i=1}^{k}\frac{1}{s_{i}^{2}+\tau^{2}}}+\tau^{2}$ (10)
implying for the specific case of $k=1$ that
$\mathrm{E}[\theta_{2}|y_{1},s_{1},\tau]$ $=$ $\mu_{\mathrm{p}}\frac{s_{1}^{2}+\tau^{2}}{\sigma_{\mathrm{p}}^{2}+s_{1}^{2}+\tau^{2}}+y_{1}\frac{\sigma_{\mathrm{p}}^{2}}{\sigma_{\mathrm{p}}^{2}+s_{1}^{2}+\tau^{2}}$ (11)
$\mathrm{Var}(\theta_{2}|y_{1},s_{1},\tau)$ $=$ $\frac{1}{\frac{1}{\sigma_{\mathrm{p}}^{2}}+\frac{1}{s_{1}^{2}+\tau^{2}}}+\tau^{2}$ (12)
(see Röver (2020) 7). One can already see that in the limiting case of an increasingly vague effect prior ($\sigma_{\mathrm{p}}\rightarrow\infty)$, the prior’s influence vanishes, and the (conditional) variance increases.
Specification of an improper uniform prior for the overall mean effect $\mu$ also leads to a proper posterior; the posterior predictive moments then are of the slightly simpler form
$\mathrm{E}[\theta_{k+1}|y_{1},\ldots,y_{k},s_{1},\ldots,s_{k},\tau]$ $=$ $\frac{\sum_{i=1}^{k}\frac{y_{i}}{s_{i}^{2}+\tau^{2}}}{\sum_{i=1}^{k}\frac{1}{s_{i}^{2}+\tau^{2}}}$ (13)
$\mathrm{Var}(\theta_{k+1}|y_{1},\ldots,y_{k},s_{1},\ldots,s_{k},\tau)$ $=$ $\frac{1}{\sum_{i=1}^{k}\frac{1}{s_{i}^{2}+\tau^{2}}}+\tau^{2}$ (14)
(see Röver (2020) 7). In the case of a single study ($k=1$), these expressions then simplify to
$\mathrm{E}[\theta_{2}|y_{1},s_{1},\tau]\;=\;y_{1}\quad\mbox{and}\quad\mathrm{Var}(\theta_{2}|y_{1},s_{1},\tau)\;=\;s_{1}^{2}+2\tau^{2}$ (15)
A.2 Gain in effective sample size
The relative gain in information from a prior may be quantified using the gain in effective sample size, which is based on the relative width of 95% intervals with and without consideration of the (informative) prior. First, consider the relative width $q$, the ratio of interval widths (or standard errors) ($q=\frac{\mbox{\small width using informative prior}}{\mbox{\small width using vague prior}}$). Assuming that standard errors are proportional to the inverse of the square root of the sample size, the gain in effective sample size then is given by $q^{-2}-1$. For example, if the informative prior yields an interval that is only half as wide ($q=0.5$), this would otherwise have required a quadrupled sample size, or a $q^{-2}-1=300\%$ increase. If the interval is 90% as wide ($q=0.9$), this corresponds to an approximate $q^{-2}-1=23\%$ increase 9.
A.3 Heterogeneity priors
Figure A7 shows several prior densities discussed in Section 3.1.2 that are all scaled to a common median (of 0.34, the median of a HN(0.5) distribution).
Figure A7: Illustration of the several heterogeneity priors compared in Section 3.1.2 in terms of their probability density functions. All priors are scaled such that they have a common median of 0.34 (the median of a half-normal(0.5) prior; dashed line)
Figure A8: Illustration of the several heterogeneity priors compared in Section 3.1.2 in terms of their cumulative distribution functions (CDFs). All priors are scaled such that they have a common median of 0.34 (dashed line)
A.4 Variance of the MAP prior
@@
Table A3: Expected values of $\tau^{2}$ based on various common (prior) distributions for $\tau$, depending on their scale parameter $s$. An asterisk ($\ast$) indicates that there is no simple analytical expression. The half-Cauchy prior would be an additional option, but does not have a finite expectation (it is in fact also a special case of the half-Student-$t$ prior, with $\nu=1$ degree of freedom)
@@
$\tau\sim\mathrm{half}\mbox{-}\mathrm{Student}\mbox{-}t_{\nu}(s)$ $\ast$ $\ast$ $2\sqrt{\frac{\nu}{\pi}}\frac{\Gamma(\frac{\nu+1}{2})}{\Gamma(\frac{\nu}{2})(\nu-1)}s$ $(\nu>1)$ $\frac{\nu}{\nu-2}s^{2}$ $(\nu>2)$
@@
Table A3 below summarizes expectations for $\tau^{2}$ for a range of common heterogeneity priors. Note also the related Table B1 in the appendix of Röver et al. (2021) 12 giving some additional details for the prior distributions shown here.
A.5 Power prior exponent’s distribution
For any fixed heterogeneity value $\tau$, the (conditional) MAP prior is equivalent to a power prior with exponent $a_{0}=\bigl(2\frac{\tau^{2}}{s_{1}^{2}}+1\bigr)^{-1}$ (see also Section 2.5) 9. Through this functional relationship, any prior density $p_{\star}(\tau)$ for the heterogeneity implies a corresponding prior for the exponent $a_{0}$ with probability density function
@@ A.5 Power prior exponent’s distribution
Figure A9 illustrates such densities for an example value of $s_{1}=0.451$ (as in the Alport example from Section 3.1).
Figure A9: Illustration of prior distributions for the power prior exponent $a_{0}$ corresponding to certain prior distributions assumed for the heterogeneity $\tau$ (and $s_{1}=0.451$)
@@
Note that (as elaborated in Section 2.5) the mapping between $\tau$ ans $a_{0}$ always depends on the “source” study’s standard error ($s_{1}$). A prior for $\tau$ may be motivated independent of $s_{1}$, while specification of a value (or distribution) for $a_{0}$ may have odd consequences when varying the source study’s size or precision ($s_{1}$).
A.6 Sensitivity analysis (Alport example)
In the Alport example application from Section 3.1, varying the half-Normal heterogeneity prior’s scale parameter affects the precision of the corresponding MAP prior, and with that, the eventual amount of borrowing from the observational data.
Figure A10: Illustration of the effect of varying the (half-normal) heterogeneity prior’s scale on the resulting RCT shrinkage estimate from Section 3.1
Figure A10 shows how the resulting shrinkage estimate for the RCT is affected. A smaller prior scale leads to more borrowing and hence a shorter shrinkage interval eventually approaching the common-effect estimate. A larger prior scale leads to less borrowing, with the shrinkage interval eventually approaching the interval based on the RCT data alone. Since a log-HR of zero always remains included, in this example there is no “tipping point” for the heterogeneity’s prior scale parameter.
A.7 Example R code
A.7.1 Alport example
@@
A.7.2 Heart failure example